OpenStudy (anonymous):

Help Please! a cube is made up of six congruent square faces . find the length of the inner diagonal (line AH) of the cube. note that AH is the hypotenuse of right triangle ADH

OpenStudy (danjs):

Do you mean, a cube with side length x and you want the length of the diagonal through the center of the cube?

OpenStudy (anonymous):

yes please

OpenStudy (danjs):

ok so first you need to find a diagonal across one of the faces of the cube.

OpenStudy (anonymous):

its 7cm

OpenStudy (danjs):



OpenStudy (danjs):

where did you get a 7 from?

OpenStudy (danjs):

so using that diagonal as the base of a triangle, the side is length x, you need to find the hypotenuse.

OpenStudy (danjs):

The hypotenuse is the diagonal through the center of the cube.

OpenStudy (anonymous):

dose that help?

OpenStudy (danjs):

Ok.. so DH is \[DH = \sqrt{7^2 + 7^2}\]

OpenStudy (danjs):

DH^2 + AD^2 = AH^2

OpenStudy (anonymous):

2?

OpenStudy (danjs):

A side length is 7, the diagonal will be longer than a side length, so 2 is not correct.

OpenStudy (anonymous):

oh okay!

OpenStudy (anonymous):

so if i round your answer to the nearest hundredth. it would be 8?

OpenStudy (danjs):

the hundredth is the second decimal place

OpenStudy (anonymous):

so 8.7? or 8.77?

OpenStudy (anonymous):

Okay that makes sense

OpenStudy (danjs):

wait, i misscalculated

OpenStudy (anonymous):

so the diagonal line is 8.77cm

OpenStudy (danjs):

no, here.. i misstyped here is the new answer

OpenStudy (anonymous):

oh okay..

OpenStudy (danjs):

DH = \[\sqrt{7^2 + 7^2} = 7\sqrt{2}\] AD = 7 so \[AH = \sqrt{7^2 + (7\sqrt{2})^2}\]

OpenStudy (danjs):

\[AH = 7\sqrt{3} \approx 12.12\]

OpenStudy (anonymous):

mmkay